# Propagation of Condensation via Neumann Localization in the Dilute Bose Gas

### 29 Mar 2026 v2

@@ Abstract
As an application, we consider the dilute Bose gas with Neumann boundary conditions. Combining the localization method with recently established free-energy bounds [4], we propagate strong condensation estimates from the Gross–Pitaevskii scale to larger boxes of side length $R\sim a(\rho a^{3})^{-\frac{3}{4}-\eta}$.
@@ Introduction
Recent advances have provided precise control of the free energy and established strong condensation estimates in Neumann boxes whose side lengths are slightly larger than the Gross–Pitaevskii scale; see in particular [4] and [8]. These results yield quantitative bounds on the number of excited particles, but only on relatively short spatial scales.

### 04 Apr 2026 v3

@@ Abstract
As an application, we consider the dilute Bose gas with Neumann boundary conditions. Combining the localization method with recently established free-energy bounds [8], we propagate strong condensation estimates obtained in [4] from the Gross–Pitaevskii scale to larger boxes of side length $R\sim a(\rho a^{3})^{-\frac{3}{4}-\eta}$.
@@ Introduction
Recent advances have provided precise control of the free energy and established strong condensation estimates in Neumann boxes whose side lengths are slightly larger than the Gross–Pitaevskii scale; see in particular [8] and an extension in [4]. These results yield quantitative bounds on the number of excited particles, but only on relatively short spatial scales.